OpenStudy (anonymous):

You have two samples of the same gas in the same size container, with the same pressure. The gas in the first container has a Kelvin temperature four times that of the gas in the other container. 1) The ratio of the number of moles of gas in the first container compared to that in the second is? 2) The ratio of the average velocity of particles in the first container compared to that in the second is? Please explain.

OpenStudy (joannablackwelder):

PV=nRT Since P, V and R are constant, when T goes up, n must go down by the same factor.

OpenStudy (joannablackwelder):

So the gas at the higher temp would have to have 4 times less moles.

OpenStudy (joannablackwelder):

1:4

OpenStudy (anonymous):

Thank you! How about the second question, can you explain that also please

OpenStudy (joannablackwelder):

No problem. For the second question, I know the root mean squared speed = sqrt(3RT/M)

OpenStudy (joannablackwelder):

So, when T1=4T2, root mean squared speed 1 = 2*root mean squared speed 2. But I'm not sure exactly how root mean squared speed relates to average velocity of particles.

OpenStudy (anonymous):

so then what would you use?

OpenStudy (anonymous):

the answer is 2:1, but i don't get how

OpenStudy (joannablackwelder):

That is what I got using the root mean squared formula. So, I guess it is the same.

OpenStudy (anonymous):

when you used sort(3RT/M) what exactly was R, T and M equal to?

OpenStudy (joannablackwelder):

R is always constant and M would be the same for the same compound. T1=4T2

OpenStudy (joannablackwelder):

So, in comparison, root mean squared speed1 ratio to root mean squared speed2 is 2:1.

OpenStudy (anonymous):

Is the constant (R) = 0.08206? and what would the number for M be?

OpenStudy (joannablackwelder):

We don't know since we don't know what the compound is, but it doesn't matter since it will be the same for both.

OpenStudy (joannablackwelder):

Another way to think of it is (I will use u for root mean squared speed) u1=sqrt(3RT1/M) and u2=Sqrt(3RT2/M)

OpenStudy (joannablackwelder):

T1=4T2, so u1=sqrt(3R*4T2/M)

OpenStudy (joannablackwelder):

Simplifying gives 2*sqrt(3RT2/M) since sqrt(4)=2

OpenStudy (joannablackwelder):

So, we can substitute to give u1=2*u2

OpenStudy (joannablackwelder):

For example, if u2 were 1, u1 would be 2. So the ratio is 2:1.

OpenStudy (anonymous):

Hm i'll have to do that on paper but thank you very much for helping out!

OpenStudy (joannablackwelder):

Yes, it is easier to see when you write it out on paper. Good luck and happy studies!